value of tan200(cot10-tan10)is

Grade:11

3 Answers

Sathya
12 years ago

The answer is 2....

Vishal Kamalakannan
12 years ago

Write tan200 as tan(180+20)=tan20.

cot10-tan10=(cos10)^2-(sin10)^2/(sin10cos10)

                 =2*cot20.

The product of the two is 2. :)

Vaishali Senad
5 years ago
Tan200°=tan(180°+20°)
Tan200° [Cot10°-tan10°]
=Tan(180°+20°) [1/tan10°-tan10°]
=Tan20° [1-tan^2 10°/tan 10°]
=Tan(10°+10°) [1-tan^2 10°/tan 10°]
=[Tan10°+Tan10°/1-tan^10°] [1- tan^2 10°/tan10°]
=2 tan 10°/tan 10°
=2.